# RovoDev Code Reviewer: A Large-Scale Online Evaluation of LLM-based Code Review Automation at Atlassian

### v1

@@ 5.1. How well do RovoDev-generated comments align with human-written comments?
Dataset. We curated our own code review benchmark dataset from Atlassian’s proprietary software repositories. However, not all code changes have associated PR and Jira information. Our initial dataset consists of 3,492 code changes and their associated high-quality, human-written code review comments. To ensure that the experiment is in a controlled setting, we filtered out the code changes that do not have pull request and Jira information. Thus, we removed 872 code changes that do not have both associated PR and Jira information. Prior studies shown that many comments written by human could be noisy (e.g., “Looks good to me”, or “/nit the space!””[liu2025too]. Following Liu et al [liu2025too], we exploited an LLM to filter out 270 code changes that are associated with noisy human-written comments (i.e., the affirmation and humour types of comments). Code review practices have matured over time, and recent comments reflect the latest standards and methodologies. To ensure that the comments are aligned with current best practices and organizational goals, we focused on the code review comments written by human reviewers over the past 3 years (2022-2025), removing 282 code changes. In summary, our internal benchmark dataset consists of 2,068 code changes that are associated 2,894 high-quality, human-written code review comments, spanning across 1,468 PRs.
@@
Evaluation Measures. We consider the following two evaluation aspects. Human-aligned Review Localization refers to the capability to generate comments at locations that closely correspond to where a human reviewer would naturally place them [hong2022commentfinder, hong2022should], while Human-aligned Comment Generation focuses on the capability to generate comments that is similar to human-written comments [hong2025retrieval, xu2020towards]. In the ideal scenario, RovoDev-generated comments should be at the same location and textually-similar as human-written comments. Therefore, we measure:
@@
%HAC (Human-Aligned Comments, capturing both location & semantic similarity) measures the percentage of human-aligned comments out of all comments generated by RovoDev. We consider a code review comment to be human-aligned if it meets two conditions: (1) it is at a close-proximity location (i.e., at exactly the same file path and within a $\pm$10 line range of the actual location); and (2) it is semantically highly similar to the human-written comment (i.e., similar comment intent). To compute the semantic similarity, we use LLM-as-a-Judge [li2024generation, zheng2023judging] ($J_{\mathrm{sim}}$) to evaluate if the set of generated code review comments ($R$) are semantically similar to a given human-written comment. Similar to Section 3.2, our LLM Judge ($J_{\mathrm{sim}}$) is based on a prompting technique using a gpt-4o-mini model where our prompt consists of persona, task definition, instructions, and assessment guidelines. We employ a score-based evaluation method [li2024generation] that assigns a similarity score to each pair of code review comments: 1 (no similarity), 2 (weak similarity), 3 (high similarity), and 4 (very high similarity). Pairs receiving a score of 3 or 4 are considered to exhibit high similarity to the human-written comment. Nevertheless, it is possible that RovoDev may generate comments that are not aligned with human. Thus, we measure %!HAC (Human-Not-Aligned Comments) as the percentage of RovoDev-generated comments that are not aligned with the human-written comments. Since RovoDev Code Reviewer is operated at the pull request level, to gain better insights, we also operationalize the following metrics at the pull request level.
@@
From the industrial perspectives, code review comments can be written in many different forms, regardless of their intended purpose. Traditional methods for evaluating these comments often depend on similarity scores (such as BLEU) or exact matches [tufano2022using, tufano2021towards, liu2025too], which restrict their ability to truly measure the quality and relevance of the generated code review comment. As a result, valuable insights that are expressed differently but share the same intent may be missed, leading to an incomplete evaluation of a given code review comment.
@@ 5.2. Which types of prompt components are most important for RovoDev Code Reviewer?
Results. Review Guideline together with Task Definition and Code Changes are the most important prompt components for code review comment generation, as it improves RovoDev’s review localization capability by 5 percentage points. Figure 5 presents the percentage impact of the prompt components on RovoDev’s effectiveness. Positive impact indicates the positive benefit that a component contributes to RovoDev (i.e., an absolute percentage loss when such the treatment component is removed from the control group). The results show that the review guideline contributes 5% positively to the RovoDev’s review localization. In other words, when the review guideline is removed from the control group, %HAC(LocationOnly) of RovoDev is decreased by 5 percentage points. Nevertheless, other types of contextual information (e.g., persona, Chain-of-Thought, PR information, and Jira Information) have a minimal impact on the RovoDev’s review localization and generation capability—the absolute percentage difference is only 1%-3%. For Jira issues and pull requests, our deeper investigation found that such information is generally high-level and lack sufficient detail to effectively generate code review comments. For persona, this finding is similar to the literature [pornprasit2024fine]—i.e., persona does not have much impact on the code review automation.
@@ 5.3. Which comment quality check components are most effective?
Results. The Actionability Check helps RovoDev to generate 15% more comments with correct locations, while the Factual Correctness Check has a minimal impact on the RovoDev’s overall effectiveness. Figure 5 shows that the ModernBERT-based Actionability Check helps improving the %HAC(LocationOnly) by 20 percentage points, and improving %HAC(Location+Semantic) by 15 percentage points. On the other hand, the Factual Correctness based on the LLM-as-a-Judge, which tends to be expensive, has a minimal impact on the RovoDev’s overall effectiveness. This finding is surprising as it contradicts the findings in the ML literature, who found that such the zero-resource black-box Judge can reduce hallucination [manakul2023selfcheckgpt, chern2023factool].
@@ 6. Threats to the Validity
Threats to construct validity. To evaluate the human alignment of our RovoDev-generated comments, we use LLM-as-a-Judge [zheng2023judging] as a proxy to measure the semantic similarity of a given code review comment to the associated human-written comment. Nevertheless, it is crucial that the Judge itself is highly reliable to ensure robust and trustworthy evaluation results [gu2024survey]. To evaluate the reliability of our Judge, we collected a small sample of 47 human-written comments. Then, we compared our Judge (gpt-4o-mini) with other models. Results show that our Judge produces scores that are highly correlated with human judgement (a Spearman correlation of 0.69), outperforming other LLMs (0.65 for mixtral-8-7b, 0.57 for llama3.1-8b, 0.54 for llmma3.0) and other cosine similarity scores (¡0.52), indicating that our Judge is of reliable enough.

### v2

@@ 5.1. How well do RovoDev-generated comments align with human-written comments?
Dataset. We curated our own code review benchmark dataset from Atlassian’s proprietary software repositories. However, not all code changes have associated PR and Jira information. Our initial dataset consists of 3,492 code changes and their associated high-quality, human-written code review comments. To ensure that the experiment is in a controlled setting, we filtered out the code changes that do not have pull request and Jira information. Thus, we removed 872 code changes that do not have both associated PR and Jira information. Prior studies shown that many comments written by human could be noisy (e.g., “Looks good to me”, or “/nit the space!””[26]. Following Liu et al [26], we exploited an LLM to filter out 270 code changes that are associated with noisy human-written comments (i.e., the affirmation and humour types of comments). Code review practices have matured over time, and recent comments reflect the latest standards and methodologies. To ensure that the comments are aligned with current best practices and organizational goals, we focused on the code review comments written by human reviewers over the past 3 years (2022-2025), removing 282 code changes. In summary, our internal benchmark dataset consists of 2,068 code changes that are associated 2,894 high-quality, human-written code review comments, spanning across 1,468 PRs.
@@
Evaluation Measures. We consider the following two evaluation aspects. Human-aligned Review Localization refers to the capability to generate comments at locations that closely correspond to where a human reviewer would naturally place them [20, 21], while Human-aligned Comment Generation focuses on the capability to generate comments that is similar to human-written comments [19, 55]. In the ideal scenario, RovoDev-generated comments should be at the same location and textually-similar as human-written comments. Therefore, we measure:
@@
%HAC (Human-Aligned Comments, capturing both location & semantic similarity) measures the percentage of human-aligned comments out of all comments generated by RovoDev. We consider a code review comment to be human-aligned if it meets two conditions: (1) it is at a close-proximity location (i.e., at exactly the same file path and within a $\pm$10 line range of the actual location); and (2) it is semantically highly similar to the human-written comment (i.e., similar comment intent). To compute the semantic similarity, we use LLM-as-a-Judge [24, 57] ($J_{\mathrm{sim}}$) to evaluate if the set of generated code review comments ($R$) are semantically similar to a given human-written comment. Similar to Section 3.2, our LLM Judge ($J_{\mathrm{sim}}$) is based on a prompting technique using a gpt-4o-mini model where our prompt consists of persona, task definition, instructions, and assessment guidelines. We employ a score-based evaluation method [24] that assigns a similarity score to each pair of code review comments: 1 (no similarity), 2 (weak similarity), 3 (high similarity), and 4 (very high similarity). Pairs receiving a score of 3 or 4 are considered to exhibit high similarity to the human-written comment. Nevertheless, it is possible that RovoDev may generate comments that are not aligned with human. Thus, we measure %!HAC (Human-Not-Aligned Comments) as the percentage of RovoDev-generated comments that are not aligned with the human-written comments. Since RovoDev Code Reviewer is operated at the pull request level, to gain better insights, we also operationalize the following metrics at the pull request level.
@@
From the industrial perspectives, code review comments can be written in many different forms, regardless of their intended purpose. Traditional methods for evaluating these comments often depend on similarity scores (such as BLEU) or exact matches [48, 49, 26], which restrict their ability to truly measure the quality and relevance of the generated code review comment. As a result, valuable insights that are expressed differently but share the same intent may be missed, leading to an incomplete evaluation of a given code review comment.
@@ 5.2. Which types of prompt components are most important for RovoDev Code Reviewer?
Results. Review Guideline together with Task Definition and Code Changes are the most important prompt components for code review comment generation, as it improves RovoDev’s review localization capability by 5 percentage points. Figure 5 presents the percentage impact of the prompt components on RovoDev’s effectiveness. Positive impact indicates the positive benefit that a component contributes to RovoDev (i.e., an absolute percentage loss when such the treatment component is removed from the control group). The results show that the review guideline contributes 5% positively to the RovoDev’s review localization. In other words, when the review guideline is removed from the control group, %HAC(LocationOnly) of RovoDev is decreased by 5 percentage points. Nevertheless, other types of contextual information (e.g., persona, Chain-of-Thought, PR information, and Jira Information) have a minimal impact on the RovoDev’s review localization and generation capability—the absolute percentage difference is only 1%-3%. For Jira issues and pull requests, our deeper investigation found that such information is generally high-level and lack sufficient detail to effectively generate code review comments. For persona, this finding is similar to the literature [34]—i.e., persona does not have much impact on the code review automation.
@@ 5.3. Which comment quality check components are most effective?
Results. The Actionability Check helps RovoDev to generate 15% more comments with correct locations, while the Factual Correctness Check has a minimal impact on the RovoDev’s overall effectiveness. Figure 5 shows that the ModernBERT-based Actionability Check helps improving the %HAC(LocationOnly) by 20 percentage points, and improving %HAC(Location+Semantic) by 15 percentage points. On the other hand, the Factual Correctness based on the LLM-as-a-Judge, which tends to be expensive, has a minimal impact on the RovoDev’s overall effectiveness. This finding is surprising as it contradicts the findings in the ML literature, who found that such the zero-resource black-box Judge can reduce hallucination [30, 7].
@@ 6. Threats to the Validity
Threats to construct validity. To evaluate the human alignment of our RovoDev-generated comments, we use LLM-as-a-Judge [57] as a proxy to measure the semantic similarity of a given code review comment to the associated human-written comment. Nevertheless, it is crucial that the Judge itself is highly reliable to ensure robust and trustworthy evaluation results [17]. To evaluate the reliability of our Judge, we collected a small sample of 47 human-written comments. Then, we compared our Judge (gpt-4o-mini) with other models. Results show that our Judge produces scores that are highly correlated with human judgement (a Spearman correlation of 0.69), outperforming other LLMs (0.65 for mixtral-8-7b, 0.57 for llama3.1-8b, 0.54 for llmma3.0) and other cosine similarity scores (¡0.52), indicating that our Judge is of reliable enough.